Suppose the real rate is 7.5 percent and the inflation rate is 3 percent. What rate would you expect to see on a Treasury bill?
Answer #1

Based on Fischer Relation,

(1 + Nominal rate) = (1 + Real Rate) * (1 + Inflation)

(1 + Nominal rate) = (1 + 7.5%) * (1 + 3%)

(1 + Nominal Rate) = 1.10725

Nominal Rate = 0.10725 = 10.725% ---> Answer

Some text mention an approximate form of Fischer Relation, which is

Nominal Rate = Real rate + Inflation

Nominal rate = 7.5% + 3% = 10.5%
Answer #2

ANSWER :



Rate on treasury bill 

= (1 + R)(1 +I) - 1 

= (1 + 0.075)(1 + 0.03) - 1

= 0.1073

= 10.73%  (ANSWER).

answered by: Tulsiram Garg